# Law of Large Numbers for continuous $N$-particle ensembles at fixed temperature

### v1

@@ 5.3 Proof of the “only if” part of the main theorem
Assume that $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ satisfies the LLN. In this section, let $\alpha=\theta^{-1}$. For each partition $\lambda$ with $|\lambda|=d$, we will show that $a^{N}_{\lambda}$ has the asymptotics as in Definition 3.2. To do this, it suffices to consider $N>\max(\alpha,1)\cdot d$, so that Theorem 5.6 applies.
We start with the case when $\lambda=(d)$. First observe that for each $d=1,2,\dots$,
$a^{N}_{(d)}=\text{the coefficient of }p_{d}(\vec{x})\ \text{in}\ \ln\mathbb{E}[B_{a}(\vec{x},\theta)]=\mathbb{E}[\ln(B_{a}(\vec{x},\theta))],$
which is therefore equal to
@@
$\frac{1}{d}\sum_{g\in\frac{1}{2}\mathbb{Z}_{\geq 0}}\sum_{(\mathbf{M},c)\in C^{(\infty)}_{d,\bullet},\mathbf{M}\ \text{normal},g(\mathbf{M})=g}\alpha^{d-1}(\alpha-1)^{\nu_{\rho}(\mathbf{M},c)}\mathbb{E}\left[\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\right]\\ \times N^{2-2g-\ell(\boldsymbol{\mu_{1}(\mathbf{M})})}\cdot f_{\eta(\mathbf{M})}\left(\left\{\left(-1\right)^{-k}\right\}_{k\geq 1}\right),$ (41)
by Theorem 5.6. Since
$\mathbb{E}\left[\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\right]\rightarrow\prod_{v\in V_{0}(\mathbf{M})}m_{deg(v)},$
as $N\rightarrow\infty$, then it is of order $O(1)$. Note that the exponent of $N$ in the sum (41) is at most $1$ and the leading terms are those with $g=0$ and $\ell(\boldsymbol{\mu_{1}}(\mathbf{M}))=1$ . The contribution to this leading term come from those constellations $\mathbf{M}$ with only one face, so that $\ell(\boldsymbol{\mu_{1}}(\mathbf{M}))=1$, and genus $g(\mathbf{M})=0$, so that $\mathbf{M}$ is embedded in a sphere and therefore is orientable. Note that $\nu_{\rho}(\mathbf{M},c)=0$ because $\mathbf{M}$ is orientable, therefore the limit
@@
$\lim_{N\rightarrow\infty}\frac{a^{N}_{(d)}}{N}=\frac{1}{d}\sum_{(\mathbf{M},c)\in C^{(\infty)}_{d,\bullet}:\ \mathbf{M}\ \text{normal},\ \ell(\boldsymbol{\mu_{1}}(\mathbf{M}))=1,\ g=0}\theta^{1-d}\prod_{v\in V_{0}(\mathbf{M})}m_{deg(v)}\cdot f_{\eta(\mathbf{M})}\left(\{(-1)^{k}\}_{k\geq 1}\right)$ (42)
exists and is finite, for all $d=1,2,\dots$.
Next, we do the case when $\ell(\lambda)\geq 2$.
Claim 5.1.
The coefficient $a^{N}_{\lambda}$ of $p_{\lambda}(x)$ in the expansion
$\ln(G_{N,\theta}(\vec{x}))=\sum_{\lambda\colon|\lambda|\leq N}{a_{\lambda}^{N}p_{\lambda}(\vec{x})}+O\big(\|x\|^{N+1}\big)$
has $N$-degree at most $\ell(\lambda)$.
Proof of the Claim.
We will deduce an expression for $a^{N}_{\lambda}$ starting from the expansion of $\ln B_{a}(\vec{;}\theta)$ in Theorem 5.6 and the following algorithm:
  • $B_{a}(\vec{x};\theta)=\exp(\ln B_{a}(\vec{x};\theta))=1+\sum_{k=1}^{\infty}\frac{1}{k!}[\ln B_{a}(\vec{x};\theta)]^{k}$,
  • $G_{N,\theta}(\vec{x})=\mathbb{E}[B_{a}(\vec{x};\theta)]=1+\sum_{k=1}^{\infty}\frac{1}{k!}\,\mathbb{E}\left([\ln B_{a}(\vec{x};\theta)]^{k}\right)$,
  • $\ln G_{N,\theta}(\vec{x})=\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j}(G_{N,\theta}(\vec{x})-1)^{j}$.
All the series above converge absolutely in a neighborhood of 0, due to Proposition-Definition 2.2. By Theorem 5.6, we have an expansion
$\ln B_{a}(\vec{x};\theta)=\sum_{\lambda}c^{N}_{\lambda}p_{\lambda}(\vec{x}),$
where each $c^{N}_{\lambda}$ is meromorphic in $N$, and is a homogeneous symmetric polynomial of degree $|\lambda|$ in $a$. If we formally treat each product of power sums of the form $\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}$ as an order $O(1)$ constant, then by Theorem 5.6, we have
$c^{N}_{\lambda}=N^{2-\ell(\boldsymbol{\mu_{1}})}(C(a,\lambda)+o(1)),$
where $C(a,\lambda)$ is calculated from the constellations with genus 0, and it contains sums of these products $\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}$.
Using the step one of the algorithm above, it follows that
$B_{a}(\vec{x};\theta)=\sum_{\lambda}b_{\lambda}^{N}p_{\lambda}(\vec{x}),$
where each $b^{N}_{\lambda}$ is a finite linear combination of terms $\prod_{j=1}^{m}c^{N}_{\kappa(j)}$ that satisfy
$\{\lambda_{1},\dots,\lambda_{\ell(\lambda)}\}=\{\kappa(1)_{1},...,\kappa(1)_{\ell(\kappa(1))}\}\sqcup\cdots\sqcup\{\kappa(m)_{1},...,\kappa(m)_{\ell(\kappa(m))}\}.$
In particular, $\ell(\lambda)=\sum_{j=1}^{m}\ell(\kappa(j))\geq m$.
Next, use step two of the algorithm to get $G_{N,\theta}(\vec{x})$: for this, simply replace $b^{N}_{\lambda}$ by $\mathbb{E}[b_{\lambda}^{N}]$, and equivalently each $\prod_{j=1}^{m}c^{N}_{\kappa(j)}$ by $\mathbb{E}\big[\prod_{j=1}^{m}c^{N}_{\kappa(j)}\big]$. From the discussions above, assuming LLN on $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$, the $N$-degree in each $\mathbb{E}\big[\prod_{j=1}^{m}c^{N}_{\kappa_{j}}\big]$ (still meromorphic in $N$) is upper bounded by
$2m-\sum_{j=1}^{m}\ell(\mu_{j})=2m-\ell(\lambda)\leq\ell(\lambda).$
Finally, from step three of the algorithm, we can write
$\ln G_{N,\theta}(\vec{x})=\sum_{\lambda}a^{N}_{\lambda}p_{\lambda}(\vec{x}),$
where each $a^{N}_{\lambda}$ is a polynomial of the $b^{N}_{\kappa}$’s. By the same argument as above, we still have that the $N$-degree in $a^{N}_{\lambda}$ is at most $\ell(\lambda)$. ∎
For the case when $\ell(\lambda)\geq 2$, we will show that $a^{N}_{\lambda}=O\big(N^{\ell(\lambda)-1}\big)$, effectively verifying the second condition of Definition 3.2. By comparing $\ln(G_{N,\theta})$ and $\mathbb{E}[\ln B_{a}(\vec{x};\theta)]$ (the latter can be obtained from $B_{a}(\vec{x};\theta)$ by switching the order of the step two and three in the algorithm), we see that for all $\lambda$ such that $\ell(\lambda)\geq 2$, the $N$-degree $\ell(\lambda)$ component in $a^{N}_{(d)}$ is not present in $\ln B_{a}(\vec{x};\theta)$. This is because without step two, which gives different ways of taking expectation on a term $p_{\lambda}(a)$ (e.g, $\mathbb{E}[p_{\lambda_{1}}(a)]\mathbb{E}[p_{\lambda_{2}}(a)]$ versus $\mathbb{E}[p_{\lambda_{1}}(a)p_{\lambda_{2}}(a)]$), that all contribute to $a^{N}_{\lambda}$, this part is exactly canceled out. With the action of step two, this part is (generically) nonzero, but as long as the LLN holds, each term of the form $N^{-\ell(\kappa)}p_{\kappa}(a/N)$, regardless of the way of taking expectation, converges to the same limit $\prod_{i=1}^{\ell(\kappa)}m_{\kappa_{i}}$. In conclusion, we have for $\lambda$ where $\ell(\lambda)\geq 2$,
$\lim_{N\rightarrow\infty}\frac{a^{N}_{\lambda}}{N^{\ell(\lambda)}}=0.$
Finally, the limit of $\frac{a^{N}_{(d)}}{N}$ is identified with $\frac{\theta^{1-d}\kappa_{d}}{d}$, by the “if” part of the theorem, and the classical fact that Eqn. (13) gives a 1-1 correspondence between $\{m_{k}\}_{k\geq 1}$ and $\{\kappa_{d}\}_{d\geq 1}$. ∎
@@
Theorem 3.4 and Eqn. (42) give the following reverse formula of Eqn. (13), in terms of constellations:
$\kappa_{d}=\sum_{(\mathbf{M},c)\in C^{(\infty)}_{d,\bullet}:\ \mathbf{M}\ \text{normal},\ \ell(\boldsymbol{\mu_{1}}(\mathbf{M}))=1,\ g=0}\prod_{v\in V_{0}(\mathbf{M})}m_{deg(v)}\cdot f_{\eta(\mathbf{M})}\left(\{(-1)^{k}\}_{k\geq 1}\right),$ (43)
for all $d=1,2,\dots$. Since constellations on closed orientable surfaces are in one-to-one correspondence with branched coverings of a sphere (see e.g. [CD22, Section 2]), this recovers a classical connection between free probability and geometry; see e.g. [BGF20] and the references therein.

### v2

@@ 5.2 An analytic expansion of the multivariate Bessel function
Remark 5.7.
An alternative direct argument showing that the sum over $g\in\frac{1}{2}\mathbb{Z}_{\geq 0}$ in (35) converges uniformly over compact subsets, as a function of $(a,\vec{x})\in\mathbb{C}^{N}\times\mathbb{C}^{N}$, as long as $N$ is very large, goes as follows.
Let $\big(M(e,m)_{\lambda,\eta}\big)$ be the matrix of base change between elementary and monomial symmetric functions, i.e.,
$m_{\eta}=\sum_{\lambda\colon|\lambda|=|\eta|}{M(e,m)_{\lambda,\eta}\cdot e_{\lambda}}.$
Then $f_{\eta}(x_{1},x_{2},\dots)=\sum_{\lambda\colon|\lambda|=|\eta|}{M(e,m)_{\lambda,\eta}\prod_{i\geq 1}{x_{\lambda_{i}}}}$. In particular,
$\bigg|f_{\eta}\Big(\big\{(-1)^{-k}\big\}_{k\geq 1}\Big)\bigg|\leq\sum_{\lambda\colon|\lambda|=|\eta|}{\Big|M(e,m)_{\lambda,\eta}\Big|}.$ (39)
By [KR06, Theorem 1], for each $\lambda$, the absolute value $\big|M(e,m)_{\lambda,\eta}\big|$ is upper bounded by the number of Lyndon word sequences over an alphabet $A$ of cardinality $|A|=4$, and with lengths adding up to $|\eta|$. Since the number of Lyndon words of length $\ell$ over $A$ is upper bounded by $4^{\ell}$, then $\big|M(e,m)_{\lambda,\eta}\big|$ is upper bounded by $4^{|\eta|}$ times the number of compositions of $|\eta|$, which is $2^{|\eta|-1}<2^{|\eta|}$, thus, $\big|M(e,m)_{\lambda,\eta}\big|\leq 8^{|\eta|}$. Plugging back into (39) gives
$\bigg|f_{\eta}\Big(\big\{(-1)^{-k}\big\}_{k\geq 1}\Big)\bigg|\leq 8^{|\eta|}\cdot p(|\eta|)\leq(8e)^{|\eta|},$
where $p(m)$ denotes the number of integer partitions of $m$. On the other hand, one can verify from [CD22, Definitions 3.6 – 3.8] that $0\leq\nu(\mathbf{M})\leq 2g(\mathbf{M})$, so $\big|(\alpha-1)^{\nu(\mathbf{M})}\big|\leq\max(\alpha,1)^{2g(\mathbf{M})}$. Finally, as observed in the proof above, the inner sum over $\mathbf{M}$ in (35) has a number of terms bounded by the number of $(2n+2g)$-constellations of size $d\leq n$. By [BD22, Theorem 3.1] and Stirling’s formula, this number can be upper bounded by $n^{4n(n+g)}$. From the previous considerations, the sum of absolute values corresponding to the sum over $g\in\frac{1}{2}\mathbb{Z}_{\geq 0}$ in (35) is upper bounded by $\sum_{g\in\frac{1}{2}\mathbb{Z}_{\geq 0}}{N^{2-2g}\cdot\max(\alpha,1)^{2g}\cdot(8e)^{2n+2g}\cdot n^{4n(n+g)}}$ times a constant independent of $g$. The last sum converges as long as $N>8e\cdot\max(\alpha,1)\cdot n^{2n}$. This verifies the Weierstrass M-test for the sum over $g\in\frac{1}{2}\mathbb{Z}_{\geq 0}$ in (35) and shows that it converges uniformly on compact subsets.
@@ 5.3 Proof of the “only if” part of the main theorem
Assume that $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ satisfies the fixed temperature LLN, meaning that
$\lim_{N\to\infty}{\mathbb{E}_{\mu_{N}}\left[\frac{1}{N^{s}}\prod_{j=1}^{s}{p^{N}_{k_{j}}(a)}\right]}=\prod_{j=1}^{s}{m_{k_{j}}},$ (40)
for all $s\in\mathbb{Z}_{\geq 1}$, $k_{1},\dots,k_{s}\in\mathbb{Z}_{\geq 1}$, and a sequence of real numbers $\{m_{k}\}_{k\geq 1}$. Assuming the expansions
$\ln\Big(\mathbb{E}_{\mu_{N}}\big[B_{a}(\vec{x};\theta)\big]\Big)=\sum_{\lambda\colon|\lambda|\leq N}{a_{\lambda}^{N}p_{\lambda}(\vec{x})}+O\big(\|x\|^{N+1}\big),$ (41)
we will prove that
$\lim_{N\to\infty}{\frac{a_{(d)}^{N}}{N}}\,\text{ exists and is finite, for all }d\in\mathbb{Z}_{\geq 1},$ (42)
and
$\lim_{N\to\infty}{\frac{a_{\lambda}^{N}}{N^{\ell(\lambda)}}}=0,\text{ whenever }\ell(\lambda)\geq 2.$ (43)
This will show that the sequence $\{\mu_{N}\}_{N\geq 1}$ is $\theta$-LLN-appropriate, and will conclude the proof of the “only if” part of the main theorem. As before, we set $\alpha:=\theta^{-1}$.
Step 1. Let us first prove (42). Let $d\in\mathbb{Z}_{\geq 1}$ be arbitrary. Assume that $N$ is large enough so that Theorem 5.6 can be applied. By definition, $a^{N}_{(d)}$ is the coefficient of $p_{d}(\vec{x})$ in the expansion of $\ln\mathbb{E}_{\mu_{N}}\big[B_{a}(\vec{x},\theta)\big]$, but this is equal to the coefficient of $p_{d}(\vec{x})$ in the expansion of $\mathbb{E}_{\mu_{N}}\big[\ln(B_{a}(\vec{x},\theta))\big]$, which by Theorem 5.6, equals
@@
$\frac{\alpha^{d-1}}{d}\sum_{g\in\frac{1}{2}\mathbb{Z}_{\geq 0}}\ \sum_{\mathbf{M}\in C^{(\infty)}_{d,\bullet}\colon\mathbf{M}\text{ normal},\,g(\mathbf{M})=g}(\alpha-1)^{\nu(\mathbf{M})}\cdot\mathbb{E}_{\mu_{N}}\left[\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\right]\\ \times N^{2-2g-\ell(\boldsymbol{\mu_{1}(\mathbf{M})})}\cdot f_{\eta(\mathbf{M})}\left(\left\{\left(-1\right)^{-k}\right\}_{k\geq 1}\right).$ (44)
Recall that $\boldsymbol{\mu_{1}}(\mathbf{M})$ is the partition with parts $\deg(v)$, $v\in V_{0}(\mathbf{M})$, and that $|\boldsymbol{\mu_{1}}(\mathbf{M})|=|\mathbf{M}|=d$, so the number of possible distinct expectation terms in the sum (44) is finite. Since each of these finitely many terms has a limit,
$\lim_{N\to\infty}{\mathbb{E}_{\mu_{N}}\left[\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\right]}=\prod_{v\in V_{0}(\mathbf{M})}m_{deg(v)},$
by assumption (40), then all the expectations inside the sum (44) are of order $O(1)$; moreover, the exponent of $N$ inside the sum is at most $1$ and is attained when $g=0$ and $\ell(\boldsymbol{\mu_{1}}(\mathbf{M}))=1$. Since the genus is $g(\mathbf{M})=0$, then $\mathbf{M}$ can be embedded in a sphere, so it is orientable, implying that $\nu(\mathbf{M})=0$. As a result,
@@
$\lim_{N\rightarrow\infty}\frac{a^{N}_{(d)}}{N}=\frac{\alpha^{d-1}}{d}\sum_{\begin{subarray}{c}\mathbf{M}\in C^{(\infty)}_{d,\bullet}:\ \mathbf{M}\ \text{normal},\\ \ell(\boldsymbol{\mu_{1}}(\mathbf{M}))=1,\ g(\mathbf{M})=0\end{subarray}}\ \prod_{v\in V_{0}(\mathbf{M})}{m_{deg(v)}}\cdot f_{\eta(\mathbf{M})}\left(\{(-1)^{k}\}_{k\geq 1}\right),$ (45)
finishing the proof of (42).
Step 2. It remains to prove (43). Let $\lambda$ be any partition with $\ell(\lambda)\geq 2$. Let $n:=|\lambda|$ and assume that $N$ is large enough so that Theorem 5.6 can be applied, i.e. we have the expansion
$\ln\Big(B_{a}(\vec{x};\theta)\Big)=\sum_{\nu:|\nu|\leq n}{b_{\nu}^{N}p_{\nu}(\vec{x})}+O\big(\|x\|^{n+1}\big),$
where the coefficients $b_{\nu}^{N}$ are polynomials in $\alpha$, $N$ and the products $\prod_{v\in V_{0}(\mathbf{M})}{\frac{1}{N}p_{\deg(v)}(a/N)}$. These products are of order $O(1)$, as $N\to\infty$, by the assumption (40). Also, the analytic expansion (35) shows that the leading asymptotic term in $\mathbb{E}_{\mu_{N}}\big[b_{\nu}^{N}\big]$ is of order $N^{2-\ell(\nu)}$ and comes from those constellations $\mathbf{M}$ with genus $g(\mathbf{M})=0$. In particular,
[…]
More generally,
$\lim_{N\to\infty}{\mathbb{E}_{\mu_{N}}\left[\prod_{i=1}^{k}\Big(N^{\ell(\nu^{(i)})-2}\cdot b_{\nu^{(i)}}^{N}\Big)\right]}=\prod_{i=1}^{k}{c_{\nu^{(i)}}},$ (47)
for any sequence of partitions $\nu^{(1)},\dots,\nu^{(k)}$ with $|\nu^{(1)}|,\cdots,|\nu^{(k)}|\leq n$.
Next, by expanding the exponential in the right hand side of
$\ln\Big(\mathbb{E}_{\mu_{N}}\big[B_{a}(\vec{x};\theta)\big]\Big)=\ln\left(\mathbb{E}_{\mu_{N}}\bigg[e^{\ln{B_{a}(\vec{x};\theta)}}\bigg]\right)=\ln\left(\mathbb{E}_{\mu_{N}}\bigg[e^{\sum_{\nu:|\nu|\leq n}{b_{\nu}^{N}p_{\nu}(\vec{x})}+O\big(\|x\|^{n+1}\big)}\bigg]\right)$ (48)
and by comparing the result with (41), we can find a formula for $a_{\lambda}^{N}$ in terms of the expected values $\Big\{\mathbb{E}\big[b_{\nu^{(1)}}^{N}\cdots b_{\nu^{(k)}}^{N}\big]\colon k\in\mathbb{Z}_{\geq 1},\,|\nu^{(1)}|+\dots+|\nu^{(k)}|\leq n\Big\}$. Such formula, together with (47), will prove the desired limit (43).
Example 5.8.
if $\lambda=(\lambda_{1},\lambda_{2})$ is a partition with $\ell(\lambda)=2$, then
$a_{(\lambda_{1},\lambda_{2})}^{N}=\mathbb{E}_{\mu_{N}}\Big[b_{(\lambda_{1},\lambda_{2})}^{N}\Big]+\mathbb{E}_{\mu_{N}}\Big[b_{(\lambda_{1})}^{N}b_{(\lambda_{2})}^{N}\Big]-\mathbb{E}_{\mu_{N}}\Big[b_{(\lambda_{1})}^{N}\Big]\cdot\mathbb{E}_{\mu_{N}}\Big[b_{(\lambda_{2})}^{N}\Big].$
Consequently, by (46), we have
$N^{-2}a_{(\lambda_{1},\lambda_{2})}^{N}=O(N^{-2})+\mathbb{E}_{\mu_{N}}\Big[\big(N^{-1}b_{(\lambda_{1})}^{N}\big)\big(N^{-1}b_{(\lambda_{2})}^{N}\big)\Big]-\mathbb{E}_{\mu_{N}}\Big[N^{-1}b_{(\lambda_{1})}^{N}\Big]\cdot\mathbb{E}_{\mu_{N}}\Big[N^{-1}b_{(\lambda_{2})}^{N}\Big],$
and this converges to $c_{1}c_{2}-c_{1}c_{2}=0$, as $N\to\infty$, by virtue of (47).
As in the previous example, when $\ell(\lambda)\geq 2$, we can express $a_{\lambda}^{N}$ as a linear combination of products of expectations
$\mathbb{E}_{\mu_{N}}\Big[b_{\mu^{(1)}}^{N}\cdots b_{\mu^{(\ell_{1})}}^{N}\Big]\cdot\mathbb{E}_{\mu_{N}}\Big[b_{\mu^{(\ell_{1}+1)}}^{N}\cdots b_{\mu^{(\ell_{1}+\ell_{2})}}^{N}\Big]\cdots\mathbb{E}_{\mu_{N}}\Big[b_{\mu^{(L-\ell_{s}+1)}}^{N}\dots b_{\mu^{(L)}}^{N}\Big],$ (49)
where $\mu^{(1)}\cup\cdots\cup\mu^{(L)}=\lambda$; note in particular that $\ell(\lambda)\geq L$, with equality if and only if each $\mu^{(i)}$ is a row partition. Because of (46)-(47), it follows that this product is of order
$O\Big(N^{\sum_{i=1}^{L}\big[(2-\ell(\mu^{(i)}))\big]}\Big).$
Since $\ell(\mu^{(i)})\geq 1$, then $\sum_{i=1}^{L}\big[(2-\ell(\mu^{(i)}))\big]\leq L\leq\ell(\lambda)$, with equality if and only if each $\mu^{(i)}$ is a row partition. This implies that the product (49) is of order $O\big(N^{\ell(\lambda)}\big)$, and if some $\mu^{(i)}$ has $\ell\big(\mu^{(i)}\big)\geq 2$, then (49) is of order $o\big(N^{\ell(\lambda)}\big)$. As a result, the limit of $N^{-\ell(\lambda)}a_{\lambda}^{N}$, as $N\to\infty$, depends on $\big\{b_{(d)}^{N}\colon d\in\mathbb{Z}_{\geq 1}\big\}$, but not on $\big\{b_{\nu}^{N}\colon\ell(\nu)\geq 2\big\}$, so we can ignore the $b_{\nu}$’s with $\ell(\nu)\geq 2$. As a result of this argument, together with equations (41) and (48), we have that
$\lim_{N\to\infty}{N^{-\ell(\lambda)}a_{\lambda}^{N}}=\lim_{N\to\infty}{N^{-\ell(\lambda)}}\big[p_{\lambda}(\vec{x})\big]\ln\mathbb{E}_{\mu_{N}}\bigg[e^{\sum_{d=1}^{n}{b_{(d)}^{N}p_{d}(\vec{x})}}\bigg].$ (50)
The latter coefficient of $p_{\lambda}(\vec{x})$ can be found by the moment-cumulant formula,
$\big[p_{\lambda}(\vec{x})\big]\ln\mathbb{E}_{\mu_{N}}\bigg[e^{\sum_{d=1}^{n}{b_{(d)}^{N}p_{d}(\vec{x})}}\bigg]=\sum_{\pi\in\mathcal{P}(\ell(\lambda))}{(-1)^{|\pi|-1}(|\pi|-1)!\cdot\prod_{B\in\pi}{\mathbb{E}_{\mu_{N}}\left[\prod_{j\in B}{b^{N}_{(j)}}\right]}},$ (51)
where, on the right hand side, $\mathcal{P}(\ell(\lambda))$ denotes the collection of set partitions $\pi$ of $\{1,2,\dots,\ell(\lambda)\}$, and $|\pi|$ denotes the number of blocks of $\pi$. Hence, by (47) applied to row partitions, and equalities (50)-(51), we deduce the desired limit:
$\lim_{N\to\infty}{N^{-\ell(\lambda)}a_{\lambda}^{N}}=\sum_{\pi\in\mathcal{P}(\ell(\lambda))}{(-1)^{|\pi|-1}(|\pi|-1)!\prod_{B\in\pi}{\prod_{j\in B}{c_{j}}}}$ $=\sum_{\pi\in\mathcal{P}(\ell(\lambda))}{(-1)^{|\pi|-1}(|\pi|-1)!}\cdot\prod_{j=1}^{\ell(\lambda)}{c_{j}}$
$=0\cdot\prod_{j=1}^{\ell(\lambda)}{c_{j}}=0.$
We remark that the equality used above to go from the first to the second line is a well-known identity: in fact, it is the moment-cumulant formula for calculating the $\ell(\lambda)$-th cumulant of the delta mass $\delta_{1}$, since this measure has moments $m_{k}=1$ and classical cumulants $\kappa_{\ell}=\mathbf{1}_{\{\ell=1\}}$.
In conclusion, we have obtained the desired $\lim_{N\rightarrow\infty}{N^{-\ell(\lambda)}a^{N}_{\lambda}}=0$, thus finishing the proof of (43), as well as the “only if” part of the main theorem.
@@
Eqn. (45) furnishes the reverse formula of Eqn. (13), in terms of constellations:
$\kappa_{d}=\sum_{\begin{subarray}{c}\mathbf{M}\in C^{(\infty)}_{d,\bullet}:\ \mathbf{M}\ \text{normal},\\ \ell(\boldsymbol{\mu_{1}}(\mathbf{M}))=1,\ g(\mathbf{M})=0\end{subarray}}\ \prod_{v\in V_{0}(\mathbf{M})}{m_{deg(v)}}\cdot f_{\eta(\mathbf{M})}\left(\{(-1)^{k}\}_{k\geq 1}\right),\text{ for all }d\in\mathbb{Z}_{\geq 1}.$ (52)
Since constellations on closed orientable surfaces are in one-to-one correspondence with branched coverings of a sphere (see e.g. [CD22, Section 2]), this recovers a classical connection between free probability and geometry; see [BGF20] and the references therein.